OpenStudy (anonymous):

Points (0,3) and (2,-5)

OpenStudy (anonymous):

write an equation

OpenStudy (jojo4eva):

u need the slope or y=mx=b form

OpenStudy (the_fizicx99):

You need to find the slope first..

OpenStudy (anonymous):

ok..

OpenStudy (jojo4eva):

the slope =-4. then the equation once again is y-y1=m(x-x1)

OpenStudy (anonymous):

hmm

OpenStudy (jojo4eva):

y-3=-4(x-0) = y=-4x+3

OpenStudy (jojo4eva):

so y=-4x+3

OpenStudy (anonymous):

this is so confuzing

OpenStudy (jojo4eva):

u still need help